# multiple directions/.style=for tree=#1, phantom, for relative level=1no edge, delay=!c.content/.pgfmath=content("!u"), before computing xy=l=0,s=0, multiple directions/.default=, grow subtree/.style=for tree=grow=#1, grow’ subtree/.style=for tree=grow’=#1

### v1

@@ 3.2 Adding self-loops
$(v,v)\in\mathbb{M}_{h}^{s}\Leftrightarrow Z_{h}^{s}(v,v)=w^{s}(v,v)<0.$ (13)
Intuitively, this corresponds to the fact that $u$ is matched to one of its neighbours if and only if the gain of forcing $u$ to be matched to one of its neighbours is positive. In other words, $u$ is matched by $\mathbb{M}_{h}$ if and only if $w^{s}(u,u)=Z^{s}(u,u)\geq 0$. However, this does not give information about who is the chosen neighbour of $u$.
We now define a new performance on the space of matchings on $\mathbb{T}^{s}$, we only consider perfect matchings $\mathbb{M}^{s}$ where every vertex has exactly one neighbour.
$\operatorname{perf}^{s}_{V}(\mathbb{T}^{s},\mathbb{M}^{s})=\mathbb{E}\left[\sum_{v\sim o}w^{s}(o,v)\mathbbm{1}_{(o,v)\in\mathbb{M}^{s}}\right]$
If $\mathbb{M}$ is a partial matching on $\mathbb{T}$, there is a natural projection of $\mathbb{M}$ into $\mathbb{M}^{s}$ a perfect matching on $\mathbb{T}^{s}$ where we simply identify previously unmatched vertices $v$ as matched to themselves through $(v,v)$. Note that this projection is invertible by just forgetting $(v,v)$. This gives a canonical way to extend pre-existing matchings on $\mathbb{T}$.
@@ 3.3 Optimality
Proposition 11.
The matching $(\mathbb{T},\mathbb{M}_{h})$ is optimal on $\mathbb{T}$. Furthermore, almost surely, any optimal matching $\mathbb{M}$ on $(\mathbb{T},Z_{h},\mathbb{M}_{h})$ shares the same set of unmatched vertices.
Before we embark on this proof, let us stress the fact that given any unimodular matching $\mathbb{M}$ on $\mathbb{T}$, Proposition 9 iii) allows to resample $\mathbb{M}_{h}$ on top of $(\mathbb{T},\mathbb{M})$ such that $(\mathbb{T},\mathbb{M},\mathbb{M}_{h})$ is unimodular, we can then make both matchings co-exist on the tree while retaining unimodularity.
@@
Lemma 12.
$(\mathbb{T}^{s},\mathbb{M}_{h}^{s})$ is a matching maximising $\operatorname{perf}^{s}_{V}$.
Proof of Lemma 12.
Fix $\mathbb{M}^{s}$ any unimodular matching on $\mathbb{T}^{s}$. And sample $(\mathbb{T}^{s},\mathbb{M}^{s},\mathbb{M}_{h}^{s})$ as in Proposition 9 iii).
@@
Figure 8: Illustration of the configuration in the proof of Proposition 9.
@@
$Z_{h}^{s}(v_{-2},{v_{-1}})=w^{s}(v_{-1},v_{0})-Z_{h}^{s}(v_{-1},v_{0}).$ (14)
Now, by definition of the $Z_{h}$, we also have
$Z_{h}^{s}(v_{-1},v_{0})=\max_{\begin{subarray}{c}y\sim v_{0}\\ y\neq v_{-1}\end{subarray}}((w^{s}(v_{0},y)-Z_{h}^{s}(v_{0},y))\geq w^{s}(v_{0},v_{1})-Z_{h}^{s}(v_{0},v_{1})$ (15)
@@
$\operatorname{perf_{V}}^{s}(\mathbb{T}^{s},\mathbb{M}^{s})-\operatorname{perf_{V}}^{s}(\mathbb{T}^{s},\mathbb{M}_{h}^{s})$
$\mathbb{E}[w^{s}(o,n(o))]-\mathbb{E}[w^{s}(o,n_{h}(o))]$
$=$ $\mathbb{E}\left[(w^{s}(v_{0},v_{1})-w^{s}(v_{0},v_{-1}))\mathbbm{1}_{A}\right]$
$\leq$ $\mathbb{E}\left[(Z_{h}^{s}(v_{0},v_{1})-Z_{h}^{s}(v_{-2},v_{-1}))\mathbbm{1}_{A}.\right]$
@@
To prove this, we will use a chain of intermediary equalities between the two events by applying the mass-transport principle. First, we need to show that the mass-transport principle still holds when $(\mathbb{T},\mathbb{M},Z_{h})$ is replaced by $(\mathbb{T}^{s},\mathbb{M}^{s},Z_{h}^{s})$. But $(\mathbb{T}^{s},\mathbb{M}^{s},Z_{h}^{s})$ is the image of $(\mathbb{T},\mathbb{M},Z_{h})$ through a deterministic bijective operator so taking expectations over either yields the same thing. Indeed for every $f$ that is a measurable function over the space of doubly pointed decorated trees, we want to show that:
$\mathbb{E}_{(T,M,Z_{h})\sim(\mathbb{T}^{s},\mathbb{M}^{s},Z_{h}^{s})}\left[\sum_{v\in V^{s}}f(T,M,Z_{h},o,v)\right]$
$=\mathbb{E}_{(T,M,Z_{h})\sim(\mathbb{T}^{s},\mathbb{M}^{s},Z_{h}^{s})}\left[\sum_{v\in V^{s}}f(T,M,Z_{h},v,o)\right].$
To this end we will write $(T,M,Z_{h})\sim s(\hat{T},\hat{M},\hat{Z_{h}})$ where $(\hat{T},\hat{M},\hat{Z_{h}})\sim(\mathbb{T},\mathbb{M},Z_{h})$, and notice $V=V^{s}$ so the required equality rewrites as:
$\mathbb{E}_{(T,M,Z_{h})\sim(\mathbb{T},\mathbb{M},Z_{h})}\left[\sum_{v\in V}f(s(T,M,Z_{h}),o,v)\right]$
$=\mathbb{E}_{(T,M,Z_{h})\sim(\mathbb{T},\mathbb{M},Z_{h})}\left[\sum_{v\in V}f(s(T,M,Z_{h}),v,o)\right].$
which is true by unimodularity of $(\mathbb{T},\mathbb{M},Z_{h})$.
@@
Let us recall the definition of $v_{k}$ so the desired equality rewrites as :
@@
Define the $(\mathbb{T}^{s},\mathbb{M}^{s},Z_{h}^{s})$ measurable function $f$ on the space of double rooted decorated trees:
@@
Applying mass-transport principle to $f$ we get:
$\mathbb{E}\left[\sum_{v\in V}f(\mathbb{T}^{s},\mathbb{M}^{s},Z_{h}^{s},o,v)\right]=\mathbb{E}\left[\sum_{v\in V}f(\mathbb{T}^{s},\mathbb{M}^{s},Z_{h}^{s},v,o)\right]$
Computing the first expectation yields:
@@
Computing the second expectation yields:
@@
Once again, applying the mass-transport principle to $f^{\prime}$ we get:
$\mathbb{E}\left[\sum_{v\in V}f^{\prime}(\mathbb{T}^{s},\mathbb{M}^{s},Z_{h}^{s},o,v)\right]=\mathbb{E}\left[\sum_{v\in V}f^{\prime}(\mathbb{T}^{s},\mathbb{M}^{s},Z_{h}^{s},v,o)\right].$
Computing the first expectation yields:
@@
Computing the second expectation yields:
@@
Proof of Proposition 11.
Let $(\mathbb{T},\mathbb{M})$ an optimal matching on $\mathbb{T}$ and $n$ its associated neighbour function. Let $(\mathbb{T},\mathbb{M},Z_{h})$ such that the law of the marginals $(\mathbb{T},\mathbb{M})$ and $(\mathbb{T},Z_{h})$ are preserved. The main idea is that by adding self-loops with the rules given in the previous subsection, the matching $\mathbb{M}_{h}^{s}$ is the most penalised perfect matching of $\mathbb{T}^{s}$. This allows us to compare $\mathbb{M}$ and $\mathbb{M}_{h}$ by comparing their counterparts on $\mathbb{T}^{s}$. The fact that $\mathbb{M}$ is optimal translates into the fact that
$\operatorname{perf}(\mathbb{M})=\mathbb{E}\left[w^{s}(o,n(o))\mathbbm{1}_{o\neq n(o)}\right]$
is maximal among unimodular perfect matchings of $\mathbb{T}^{s}$.
Let us decompose $w^{s}(o,n(o))$ and $w^{s}(o,n_{h}(o))$ into their positive and negative parts.
$\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})$ $=\mathbb{E}\left[w^{s}(o,n(o))\mathbbm{1}_{w^{s}(o,n(o))\geq 0}\right]+\mathbb{E}\left[w^{s}(o,n(o))\mathbbm{1}_{w^{s}(o,n(o))<0}\right]$
$\operatorname{perf_{V}}^{s}(\mathbb{M}_{h}^{s})$ $=\mathbb{E}\left[w^{s}(o,n_{h}(o))\mathbbm{1}_{w^{s}(o,n_{h}(o))\geq 0}\right]+\mathbb{E}\left[w^{s}(o,n_{h}(o))\mathbbm{1}_{w^{s}(o,n_{h}(o))<0}\right]$
@@
This implies that the previous decomposition can be written as:
$\operatorname{perf_{V}}^{s}(\mathbb{M}_{h}^{s})=\operatorname{perf_{V}}(\mathbb{M}_{h})+\mathbb{E}[w^{s}(o,o)\mathbbm{1}_{w^{s}(o,o)<0}.]$
Now we can lower bound the positive part of $\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})$ by $\operatorname{perf_{V}}(\mathbb{M})$ by omitting the positive weight self-loops and we can lower bound the negative part by $[w^{s}(o,o)\mathbbm{1}_{w^{s}(o,o)<0}.]$ because $\mathbb{M}^{s}$ can at most include every negative self-loop:
$\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})\geq\operatorname{perf_{V}}(\mathbb{M})+\mathbb{E}[w^{s}(o,o))\mathbbm{1}_{w^{s}(o,o)<0}].$
@@
$\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})\geq\operatorname{perf_{V}}(\mathbb{M}_{h})+\mathbb{E}[w^{s}(o,o)\mathbbm{1}_{w^{s}(o,o)<0})]=\operatorname{perf_{V}}^{s}(\mathbb{M}_{h}^{s}).$
But we proved in Lemma 12 that $\mathbb{M}_{h}^{s}$ is optimal for $\operatorname{perf_{V}}^{s}$ so
@@
$\operatorname{perf_{V}}(\mathbb{\mathbb{M}})=\mathbb{E}\left[w^{s}(o,n(o))\mathbbm{1}_{w^{s}(o,n(o))\geq 0}\right],$
$\mathbb{E}\left[w^{s}(o,n(o))\mathbbm{1}_{w^{s}(o,n(o))<0}\right]=\mathbb{E}[w^{s}(o,o)\mathbbm{1}_{w^{s}(o,o)<0})].$
So we’ve shown that $\mathbb{M}^{s}$ maximizes $\operatorname{perf_{V}}^{s}$ and that $\mathbb{M}_{h}$ is optimal on $\mathbb{T}$. Furthermore, by unimodularity, we get that the set of unmatched vertices by $\mathbb{M}$ is almost surely the same as the set of unmatched vertices by $\mathbb{M}_{h}$, and it is almost surely the set of vertices $v$ such that $w^{s}(v,v)<0$. ∎
@@ 3.4 Uniqueness
We will keep using the vertex-rooted point of view in this section.
The goal of this section is to prove that $\mathbb{M}_{h}$ is also the unique optimal matching in law. We will actually prove the following stronger result:
Proposition 13.
Let $(\mathbb{T},\mathbb{M})$ be a unimodular optimal matched graph, sample $(\mathbb{T},\mathbb{M},Z_{h})$ with Proposition 9. Almost surely, $\mathbb{M}=\mathbb{M}_{h}$.
@@
$\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})-\operatorname{perf_{V}}^{s}(\mathbb{M}_{h}^{s})$ $=\mathbb{E}\left[(w^{s}(v_{0},v_{1})-Z_{h}^{s}(v_{0},v_{1})-Z_{h}^{s}(v_{-1},v_{0}))\mathbbm{1}_{v_{1}\neq v_{-1}}\right]$
Equality between $\operatorname{perf_{V}}^{s}(\mathbb{M}_{h}^{s})$ and $\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})$ (see Remark Remark) implies that the RHS is equal to 00 , furthermore Equation (15) says that the variable in the expectation of the RHS is non negative almost surely, we conclude that almost surely:
$Z_{h}^{s}(v_{-1},o)=\max_{\begin{subarray}{c}y\sim o\\ y\neq v_{-1}\end{subarray}}(w^{s}(o,y)-Z_{h}^{s}(o,y))=w^{s}((o,v_{1})-Z_{h}^{s}(o,v_{1})).$
@@
$v_{-1}=n_{h}(o)=\operatorname*{argmax}_{y\sim o}(w^{s}(o,y)-Z_{h}^{s}(o,y)).$
So $v_{1}$ achieves the maximum among the list of $(w^{s}(o,y)-Z_{h}^{s}(o,y))$ stripped of its maximum. We deduce that $v_{1}=n(o)$ achieves the second largest among the $(w^{s}(o,y)-Z_{h}^{s}(o,y))$ that we will write as $\overset{[2]}{\underset{y\sim o}{\operatorname*{argmax}}}(w^{s}(o,y)-Z_{h}^{s}(o,y))$.
@@
$\mathbb{P}\left(n(o)=\underset{y\sim o}{\operatorname*{argmax}}(w^{s}(o,y)-Z_{h}^{s}(o,y))\text{ or }\underset{y\sim o}{\operatorname*{\overset{[2]}{argmax}}}(w^{s}(o,y)-Z_{h}^{s}(o,y))\right)=1.$
@@
$v_{1}$ $=\underset{y\sim o}{\operatorname*{\overset{[2]}{argmax}}}(w^{s}(o,y)-Z_{h}^{s}(o,y)),$
$v_{-1}$ $=\underset{y\sim o}{\operatorname*{argmax}}(w^{s}(o,y)-Z_{h}^{s}(o,y)),$
$v_{-2k}$ $=\underset{y\sim w_{-2k+1}}{\operatorname*{\overset{[2]}{argmax}}}(w^{s}(v_{-2k+1},y)-Z_{h}^{s}(v_{-2k+1},y))$ $\forall k\in\mathbb{N}^{*},$
$v_{-2k-1}$ $=\underset{y\sim w_{-2k}}{\operatorname*{argmax}}(w^{s}(v_{-2k},y)-Z_{h}^{s}(v_{-2k},y))$ $\forall k\in\mathbb{N}^{*},$
$v_{2k}$ $=\underset{y\sim w_{2k-1}}{\operatorname*{argmax}}(w^{s}(v_{2k-1},y)-Z_{h}^{s}(v_{2k-1},y))$ $\forall k\in\mathbb{N}^{*},$
$v_{2k+1}$ $=\underset{y\sim w_{2k}}{\operatorname*{\overset{[2]}{argmax}}}(w^{s}(v_{2k},y)-Z_{h}^{s}(v_{2k},y))$ $\forall k\in\mathbb{N}^{*}.$
@@
Furthermore, by the second part of Proposition 11, almost surely, for any $v\in V$, $n_{h}(v)=v$ if and only if $n(v)=v$, from this we deduce that on $A$ we have $v_{-1}\neq v_{0}$ and, by induction, $\forall k\in\mathbb{N},v_{-k}\neq v_{-k-1}$ (no loops in the path) and also that $\forall k\in\mathbb{N},v_{-k-2}\neq v_{-k}$ (the path cannot go back up).
@@
We have that $\mathbb{P}(A)>0$ and $\mathbb{P}(C|A)=1$ so $\mathbb{P}(C)\geq\mathbb{P}(A)>0$. Furthermore, $C\subset C_{-\infty,0}$ so $\mathbb{P}(C_{-\infty,0})>0$. We are going to show that this is not possible by establishing a contradiction for the value of the probability
@@
As long as $\deg(o)\geq 1$, there is one regular edge with $o$ as an endpoint and the self-loop $\{o,o\}$ the $\operatorname*{argmax}$ and $\operatorname*{\overset{[2]}{argmax}}$ around $o$ are thus automatically different. We can then decompose $C_{-\infty,2H}$ as $C_{-\infty,0}\cap C_{0,2}\cap\{\deg(o)\geq 1\}$

### v2

@@ 3.2 Adding self-loops
$(v,v)\in\mathbb{M}_{h}^{s}\Leftrightarrow Z_{h}^{s}(v,v)=w^{s}_{h}(v,v)<0.$ (15)
Intuitively, this corresponds to the fact that $u$ is matched to one of its neighbours if and only if the gain of forcing $u$ to be matched to one of its neighbours is positive. In other words, $u$ is matched by $\mathbb{M}_{h}$ if and only if $w^{s}_{h}(u,u)=Z^{s}(u,u)\geq 0$. However, this does not give information about who is the chosen neighbour of $u$.
We will later use the fact that introducing self loops with this specific choice of weights conserves unimodularity:
Lemma 3.4.
Let $(\mathbb{T},(f_{i})_{i\in\{0,...,I\}})$ be a decorated weighted (vertex rooted) UBGW random tree. Fix $h$ a solution to Equation (10) and recall the definition of $\mathbb{T}^{s}_{h}$ in Equation (12). Then $(\mathbb{T}^{s}_{h},Z_{h}^{s},(f_{i}^{s})_{i\in\{0,...,I\}})$ is unimodular.
Proof.
We need to show that the mass-transport principle still holds for $(\mathbb{T}^{s}_{h},(f_{i}^{s})_{i\in\{0,...,I\}})$. The idea is that $(\mathbb{T}^{s}_{h},(f_{i}^{s})_{i\in\{0,...,I\}})$ is the image of $(\mathbb{T},Z_{h},f_{i})$ through a deterministic bijective operator that doesn’t modify vertices. For every $f$ that is a measurable function over the space of doubly pointed decorated trees, we want to show that:
$\mathbb{E}_{(T,Z,f_{i})\sim(\mathbb{T}^{s}_{h},Z_{h}^{s},f_{i}^{s})}\left[\sum_{v\in V^{s}}f(T,Z_{h},(f_{i}),o,v)\right]=\mathbb{E}_{(T,Z,f_{i})\sim(\mathbb{T}^{s}_{h},Z_{h}^{s},f_{i}^{s})}\left[\sum_{v\in V^{s}}f(T,Z_{h},(f_{i}),v,o)\right].$
To this end we will write $(T^{s},Z_{h}^{s},f_{i}^{s})\sim s(\hat{T},Z_{h},f_{i}))$ where $(\hat{T},\hat{Z_{h}},(\hat{f_{i}^{s}}))\sim(\mathbb{T},Z_{h})$, and notice that neither the vertex set nor the root have changed so the required equality rewrites as:
$\mathbb{E}_{(T,Z_{h},f_{i})\sim(\mathbb{T},Z_{h},(f_{i})}\left[\sum_{v\in V}f(s(T,Z_{h},(f_{i})),o,v)\right]=\mathbb{E}_{(T,Z_{h},f_{i})\sim(\mathbb{T},Z_{h},(f_{i})}\left[\sum_{v\in V}f(s(T,Z_{h},(f_{i})),v,o)\right].$
which is true by unimodularity of $(\mathbb{T},Z_{h},(f_{i}))$. ∎
Finally, we define a new performance on the space of matchings on $\mathbb{T}^{s}_{h}$, we only consider perfect matchings $\mathbb{M}^{s}$ where every vertex has exactly one neighbour.
$\operatorname{perf}^{s}_{V}(\mathbb{T}^{s}_{h},\mathbb{M}^{s})=\mathbb{E}\left[\sum_{v\sim o}w^{s}_{h}(o,v)\mathbbm{1}_{(o,v)\in\mathbb{M}^{s}}\right].$ (16)
The aim of the next section is to first prove that $\mathbb{M}_{h}^{s}$ always maximises $\operatorname{perf}^{s}_{V}$ among unimodular perfect matchings of $(\mathbb{T}^{s}_{h},w^{s}_{h})$. As a consequence, as can be guessed from (15), we will then show that $\mathbb{M}_{h}$ is optimal among unimodular matchings of $(\mathbb{T},w)$.
@@ 3.3 Optimality
Proposition 3.5.
For any solution $h$ to Equation (10), the matching $(\mathbb{T},\mathbb{M}_{h})$ is optimal on $\mathbb{T}$. Furthermore, almost surely, any unimodular optimal matching $\mathbb{M}$ on $(\mathbb{T},Z_{h},\mathbb{M}_{h})$ shares the same set of unmatched vertices.
Before stating the proof, let us stress the fact that given any unimodular matching $\mathbb{M}$ on $\mathbb{T}$, Proposition 3.2 ii) states that $\mathbb{M}_{h}$ can be resampled on top of $(\mathbb{T},\mathbb{M})$ such that $(\mathbb{T},\mathbb{M},\mathbb{M}_{h})$ is unimodular, and both matchings co-exist on the tree while retaining unimodularity.
@@
Lemma 3.6.
Fix $h$ a solution to Equation (10) and recall the definition of $\mathbb{T}^{s}_{h}$ in Equation (12). Given $(\mathbb{T},Z_{h})$, the matching $\mathbb{M}_{h}^{s}$ maximises $\operatorname{perf}^{s}_{V}$ among all unimodular perfect matchings of $\mathbb{T}^{s}_{h}$.
Proof of Lemma 3.6.
Fix $\mathbb{M}^{s}$ any unimodular matching on $\mathbb{T}^{s}_{h}$. And sample $(\mathbb{T}^{s}_{h},\mathbb{M}^{s},\mathbb{M}_{h}^{s})$ as in Proposition 3.2 ii).
@@
Figure 8: Illustration of the configuration in the proof of Proposition 3.2.
@@
$Z_{h}^{s}(v_{-2},{v_{-1}})=w^{s}_{h}(v_{-1},v_{0})-Z_{h}^{s}(v_{-1},v_{0}).$ (17)
Now, by definition of the $Z_{h}^{s}$, we also have
$Z_{h}^{s}(v_{-1},v_{0})=\max_{\begin{subarray}{c}y\sim v_{0}\\ y\neq v_{-1}\end{subarray}}((w^{s}_{h}(v_{0},y)-Z_{h}^{s}(v_{0},y))\geq w^{s}_{h}(v_{0},v_{1})-Z_{h}^{s}(v_{0},v_{1})$ (18)
@@
$\operatorname{perf_{V}}^{s}(\mathbb{T}^{s}_{h},\mathbb{M}^{s})-\operatorname{perf_{V}}^{s}(\mathbb{T}^{s}_{h},\mathbb{M}_{h}^{s})$ $=\mathbb{E}[w^{s}_{h}(o,n(o))]-\mathbb{E}[w^{s}_{h}(o,n_{h}(o))]$
$=\mathbb{E}\left[(w^{s}_{h}(v_{0},v_{1})-w^{s}_{h}(v_{0},v_{-1}))\mathbbm{1}_{A}\right]$
$\leq\mathbb{E}\left[(Z_{h}^{s}(v_{0},v_{1})-Z_{h}^{s}(v_{-2},v_{-1}))\mathbbm{1}_{A}\right].$
@@
To prove this, we will use a chain of intermediary equalities between the two events by applying the mass-transport principle, which is possible thanks to Lemma 3.4.
@@
Let us recall the definition of $v_{k}$ so the desired equality rewrites as
@@
Define the measurable function $f$ on the space of double rooted decorated trees
@@
Applying mass-transport principle to $f$ we get
$\mathbb{E}\left[\sum_{v\in V}f(\mathbb{T}^{s}_{h},\mathbb{M}^{s},Z_{h}^{s},o,v)\right]=\mathbb{E}\left[\sum_{v\in V}f(\mathbb{T}^{s}_{h},\mathbb{M}^{s},Z_{h}^{s},v,o)\right]$
Computing the first expectation yields
@@
Computing the second expectation yields
@@
Once again, applying the mass-transport principle to $f^{\prime}$ we get
$\mathbb{E}\left[\sum_{v\in V}f^{\prime}(\mathbb{T}^{s}_{h},\mathbb{M}^{s},Z_{h}^{s},o,v)\right]=\mathbb{E}\left[\sum_{v\in V}f^{\prime}(\mathbb{T}^{s}_{h},\mathbb{M}^{s},Z_{h}^{s},v,o)\right].$
Computing the first expectation yields
@@
Computing the second expectation yields
@@
Proof of Proposition 3.5.
Let $(\mathbb{T},\mathbb{M})$ an optimal matching on $\mathbb{T}$ and $n$ its associated neighbour function. Let $(\mathbb{T},\mathbb{M},Z_{h})$ such that the law of the marginals $(\mathbb{T},\mathbb{M})$ and $(\mathbb{T},Z_{h})$ are preserved. The main idea is that by adding self-loops with the rules given in the previous subsection, the matching $\mathbb{M}_{h}^{s}$ is the most penalised perfect matching of $\mathbb{T}^{s}_{h}$. This allows us to compare $\mathbb{M}$ and $\mathbb{M}_{h}$ by comparing their counterparts on $\mathbb{T}^{s}_{h}$. The fact that $\mathbb{M}$ is optimal translates into the fact that
$\operatorname{perf_{V}}(\mathbb{M})=\mathbb{E}\left[w^{s}_{h}(o,n(o))\mathbbm{1}_{o\neq n(o)}\right]$
is maximal among unimodular perfect matchings of $\mathbb{T}^{s}_{h}$.
Let us decompose $w^{s}_{h}(o,n(o))$ and $w^{s}_{h}(o,n_{h}(o))$ into their positive and negative parts. We denote by $\mathbb{M}^{s}$ the perfect matching of $\mathbb{T}_{h}^{s}$ corresponding to $\mathbb{M}$.
$\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})$ $=\mathbb{E}\left[w^{s}_{h}(o,n(o))\mathbbm{1}_{w^{s}_{h}(o,n(o))\geq 0}\right]+\mathbb{E}\left[w^{s}_{h}(o,n(o))\mathbbm{1}_{w^{s}_{h}(o,n(o))<0}\right]$
$\operatorname{perf_{V}}^{s}(\mathbb{M}_{h}^{s})$ $=\mathbb{E}\left[w^{s}_{h}(o,n_{h}(o))\mathbbm{1}_{w^{s}_{h}(o,n_{h}(o))\geq 0}\right]+\mathbb{E}\left[w^{s}_{h}(o,n_{h}(o))\mathbbm{1}_{w^{s}_{h}(o,n_{h}(o))<0}\right].$
@@
This implies that the previous decomposition can be written as
$\operatorname{perf_{V}}^{s}(\mathbb{M}_{h}^{s})=\operatorname{perf_{V}}(\mathbb{M}_{h})+\mathbb{E}[w^{s}_{h}(o,o)\mathbbm{1}_{w^{s}_{h}(o,o)<0}].$
Now we can lower bound the positive part of $\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})$ by $\operatorname{perf_{V}}(\mathbb{M})$ by omitting the positive weight self-loops and we can lower bound the negative part by $w^{s}_{h}(o,o)\mathbbm{1}_{w^{s}_{h}(o,o)<0}.$ because $\mathbb{M}^{s}$ can at most include every negative self-loop:
$\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})\geq\operatorname{perf_{V}}(\mathbb{M})+\mathbb{E}[w^{s}_{h}(o,o))\mathbbm{1}_{w^{s}_{h}(o,o)<0}].$
@@
$\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})\geq\operatorname{perf_{V}}(\mathbb{M}_{h})+\mathbb{E}[w^{s}_{h}(o,o)\mathbbm{1}_{w^{s}_{h}(o,o)<0})]=\operatorname{perf_{V}}^{s}(\mathbb{M}_{h}^{s}).$
However, we proved in Lemma 3.6 that $\mathbb{M}_{h}^{s}$ is optimal for $\operatorname{perf_{V}}^{s}$ so
@@
$\operatorname{perf_{V}}(\mathbb{\mathbb{M}})=\mathbb{E}\left[w^{s}_{h}(o,n(o))\mathbbm{1}_{w^{s}_{h}(o,n(o))\geq 0}\right],$
$\mathbb{E}\left[w^{s}_{h}(o,n(o))\mathbbm{1}_{w^{s}_{h}(o,n(o))<0}\right]=\mathbb{E}[w^{s}_{h}(o,o)\mathbbm{1}_{w^{s}_{h}(o,o)<0})].$
So we have shown that $\mathbb{M}^{s}$ maximizes $\operatorname{perf_{V}}^{s}$ and that $\mathbb{M}_{h}$ is optimal on $\mathbb{T}$. Furthermore, by unimodularity, we get that the set of unmatched vertices by $\mathbb{M}$ is almost surely the same as the set of unmatched vertices by $\mathbb{M}_{h}$, and it is almost surely the set of vertices $v$ such that $w^{s}_{h}(v,v)<0$. ∎
@@ 3.4 Uniqueness
We will keep using the vertex-rooted point of view in this section. The goal of this section is to prove that $\mathbb{M}_{h}$ is also the unique optimal matching in law. We will actually prove the following stronger result:
Proposition 3.7.
Let $(\mathbb{T},\mathbb{M})$ be a unimodular optimal matched UBGW tree, sample $(\mathbb{T},\mathbb{M},Z_{h})$ with Proposition 3.2. Almost surely, $\mathbb{M}=\mathbb{M}_{h}$.
@@
$\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})-\operatorname{perf_{V}}^{s}(\mathbb{M}_{h}^{s})$ $=\mathbb{E}\left[(w^{s}_{h}(v_{0},v_{1})-Z_{h}^{s}(v_{0},v_{1})-Z_{h}^{s}(v_{-1},v_{0}))\mathbbm{1}_{v_{1}\neq v_{-1}}\right]$
Equality between $\operatorname{perf_{V}}^{s}(\mathbb{M}_{h}^{s})$ and $\operatorname{perf_{V}}^{s}(\mathbb{M}^{s})$ (see Remark Remark) implies that the right-hand side is equal to 00. Furthermore, Equation (18) says that the variable in the expectation of the right-hand side is non negative almost surely. We conclude that, almost surely,
$Z_{h}^{s}(v_{-1},o)=\max_{\begin{subarray}{c}y\sim o\\ y\neq v_{-1}\end{subarray}}(w^{s}_{h}(o,y)-Z_{h}^{s}(o,y))=w^{s}_{h}((o,v_{1})-Z_{h}^{s}(o,v_{1})).$
@@
$v_{-1}=n_{h}(o)=\operatorname*{argmax}_{y\sim o}(w^{s}_{h}(o,y)-Z_{h}^{s}(o,y)).$
So $v_{1}$ achieves the maximum among the list of $(w^{s}_{h}(o,y)-Z_{h}^{s}(o,y))$ stripped of its maximum. We deduce that $v_{1}=n(o)$ achieves the second largest among the $(w^{s}_{h}(o,y)-Z_{h}^{s}(o,y))$ that we will write as $\overset{[2]}{\underset{y\sim o}{\operatorname*{argmax}}}(w^{s}_{h}(o,y)-Z_{h}^{s}(o,y))$.
@@
$\mathbb{P}\left(n(o)=\underset{y\sim o}{\operatorname*{argmax}}(w^{s}_{h}(o,y)-Z_{h}^{s}(o,y))\text{ or }\underset{y\sim o}{\operatorname*{\overset{[2]}{argmax}}}(w^{s}_{h}(o,y)-Z_{h}^{s}(o,y))\right)=1.$
@@
$v_{1}$ $=\underset{y\sim o}{\operatorname*{\overset{[2]}{argmax}}}(w^{s}_{h}(o,y)-Z_{h}^{s}(o,y)),$
$v_{-1}$ $=\underset{y\sim o}{\operatorname*{argmax}}(w^{s}_{h}(o,y)-Z_{h}^{s}(o,y)),$
$v_{-2k}$ $=\underset{y\sim w_{-2k+1}}{\operatorname*{\overset{[2]}{argmax}}}(w^{s}_{h}(v_{-2k+1},y)-Z_{h}^{s}(v_{-2k+1},y))$ $\forall k\in\mathbb{N}^{\star},$
$v_{-2k-1}$ $=\underset{y\sim w_{-2k}}{\operatorname*{argmax}}(w^{s}_{h}(v_{-2k},y)-Z_{h}^{s}(v_{-2k},y))$ $\forall k\in\mathbb{N}^{\star},$
$v_{2k}$ $=\underset{y\sim w_{2k-1}}{\operatorname*{argmax}}(w^{s}_{h}(v_{2k-1},y)-Z_{h}^{s}(v_{2k-1},y))$ $\forall k\in\mathbb{N}^{\star},$
$v_{2k+1}$ $=\underset{y\sim w_{2k}}{\operatorname*{\overset{[2]}{argmax}}}(w^{s}_{h}(v_{2k},y)-Z_{h}^{s}(v_{2k},y))$ $\forall k\in\mathbb{N}^{\star}.$
@@
Furthermore, by the second part of Proposition 3.5, almost surely, for any $v\in V$, $n_{h}(v)=v$ if and only if $n(v)=v$, from this we deduce that on $A$ we have $v_{-1}\neq v_{0}$ and, by induction, $\forall k\in\mathbb{N},v_{-k}\neq v_{-k-1}$ (no loops in the path) and also that $\forall k\in\mathbb{N},v_{-k-2}\neq v_{-k}$ (the path cannot go back up).
@@
We have that $\mathbb{P}(A)>0$ and $\mathbb{P}(C|A)=1$ so $\mathbb{P}(C)\geq\mathbb{P}(A)>0$. Furthermore, $C\subset C_{-\infty,0}$ so $\mathbb{P}(C_{-\infty,0})>0$. We are going to show that this is not possible by establishing a contradiction for the value on the probability
@@
As long as $\deg(o)\geq 1$, there is one regular edge with $o$ as an endpoint and the self-loop $\{o,o\}$ the $\operatorname*{argmax}$ and $\operatorname*{\overset{[2]}{argmax}}$ around $o$ are thus automatically different. We can then decompose $C_{-\infty,2}$ as $C_{-\infty,0}\cap C_{0,2}\cap\{\deg(o)\geq 1\}$